Find the 25th, 50th, and 75th percentile from the following list of 26 data

6 8 9 20 24
30 31 42 43 50
60 62 63 70 75
77 80 83 84 86
88 89 91 92 94
99



25th percentile = 50th percentile = 75th percentile =
Answer #1

Given data is:

6,8,9,20,24,30,31,42,43,50,60,62,63,70,75,77,80,83,84,86,88,89,91,92,94,99

There are total n=26 observations.

Therefore, 25th percentile is: (n+1)/4=27/4= 6.75th observation.

That is 7th observation.

Hence, 25th percentile =31

The 50th percentile is:(n+1)/2=27/2=13.5th observation.

Hence 50th percentile=(63+70)/2=66.5

The 75th percentile is= observation.

That is 20th observation.

Hence 75th percentile =86